OpenStudy (anonymous):

For what values of b are the given vectors orthogonal? (Enter your answers as a comma-separated list.) (−4, b, 3), (b, b^2, b)

jimthompson5910 (jim_thompson5910):

The two vectors are orthogonal or perpendicular if and only if the dot product of the two vectors is 0

jimthompson5910 (jim_thompson5910):

so this means (−4, b, 3) dot (b, b^2, b) = 0 -4*b + b*b^2 + 3*b = 0 -4b + b^3 + 3b = 0 solve for b